# Towards Scene Graph Anticipation

### v1

@@ 6.1 Baselines
We select two methods from VidSGG literature, STTran [7] and DSGDetr [9], as our strong baselines [^5] Although VidSGG literature witnessed many approaches recently, most use the representation processing pipeline proposed by the two selected transformer-based VidSGG methods.for adaptations. In our adaptation of selected methods to anticipate the relationships, we retain the object representation processing unit (ORPU) and spatial context processing unit (SCPU) as proposed originally and introduce two variants of latent dynamics processing unit (LDPU) (see Fig. 3). The proposed two variants differ in (1) architecture and (2) loss functions.
Figure 3: Overview of Baselines. In our proposed Variant 1(shown to the left), we input relationship representations to an anticipatory transformer to generate relationship representations for future frames auto-regressively. A predicate classification network is then employed to decode these anticipated relationship representations. Meanwhile, in Variant 2(shown to the right), we enhance relationship representations by passing them through a temporal encoder. These representations are fed to an auto-regressive anticipatory transformer to anticipate future relationship representations. Here, we employ two predicate classification heads: one decodes the observed relationship representations, and the other decodes the anticipated relationship representations. In both variants, the auto-regressive anticipatory transformer acts as the generative model, predicting the evolution of relationships between the objects.
In terms of architectural differences, in our baseline+ variants, we employ an anticipatory transformer built using the vanilla transformer architecture [35] to generate future relationship representations by processing the relationship representations in the observed temporal context. Meanwhile, in our baseline++ variants, we introduce an additional temporal encoder to process the representations before we feed them through the anticipatory transformer. This component further refines the representations by enhancing the model’s understanding of spatio-temporal dynamics. In the context of loss functions, the approach followed by our baseline+ variants focuses solely on decoding the anticipated relationship representations. However, in contrast, our baseline++ variants adopt a more comprehensive strategy. These not only decode the anticipated representations but also simultaneously decode the observed representations. This dual-decoding approach allows for a more nuanced understanding of relationships.
@@
0.3 STTran+ [7] 12.5 19.6 20.7 13.9 21.6 40.8 3.4 5.3 5.7 3.5 7.3 20.3
DSGDetr+ [9] 12.8 19.6 20.4 14.3 21.8 41.3 3.5 5.3 5.6 3.6 7.6 21.2
STTran++ [7] 18.5 27.9 29.5 15.4 27.2 48.6 5.9 10.4 11.3 6.2 14.1 31.2
DSGDetr++ [9] 19.5 28.3 29.4 16.8 29.0 48.9 6.0 10.3 11.0 8.4 16.7 32.3
@@
0.5 STTran+ [7] 13.1 21.1 22.2 14.9 22.6 42.9 3.6 5.8 6.2 3.7 7.6 21.4
DSGDetr+ [9] 13.6 20.9 21.9 15.2 23.1 43.3 3.8 5.8 6.1 3.9 8.0 22.2
STTran++ [7] 19.7 30.2 31.8 16.6 29.1 51.5 6.3 11.3 12.3 6.6 14.7 33.4
DSGDetr++ [9] 20.7 30.3 31.6 17.4 30.5 51.9 6.4 11.0 11.7 8.4 17.0 33.9
@@
0.7 STTran+ [7] 14.9 23.4 24.7 16.6 25.1 47.2 4.1 6.5 7.0 4.2 8.5 24.0
DSGDetr+ [9] 15.5 23.4 24.3 16.8 25.3 47.4 4.3 6.5 6.9 4.3 8.8 24.7
STTran++ [7] 22.1 33.6 35.2 19.0 32.8 56.8 7.0 12.6 13.6 7.7 17.1 36.8
DSGDetr++ [9] 22.9 33.6 34.9 19.8 34.1 56.7 7.1 12.6 13.3 9.5 19.2 37.2
@@
0.9 STTran+ [7] 15.2 24.1 25.4 17.5 26.8 49.6 4.5 7.0 7.5 4.6 9.2 24.3
DSGDetr+ [9] 16.2 24.8 25.9 17.9 27.7 51.4 4.8 7.2 7.6 4.7 9.7 25.9
STTran++ [7] 23.6 35.5 37.4 20.2 35.0 60.2 7.4 13.4 14.6 8.9 18.4 38.8
DSGDetr++ [9] 24.4 36.1 37.6 22.2 37.1 61.0 7.4 13.8 14.8 11.4 21.0 39.5
@@
6.2 Results
Action Genome Scenes. We present the results in Table. 1. In this setting, we observe that both STTran++ and DSGDetr++ consistently outperform their basic counterparts. Furthermore, it can be clearly observed that the SceneSayerSDE model consistently outperforms the SceneSayerODE model, which, in turn, performs better than the baseline variants. SceneSayerODE/SDE models only fall short on the R@50 metric in the No Constraint graph generation strategy. Nevertheless, the R@K metric holds greater importance for lower values of K and, in R@10 and R@20 metrics, consistently outperforms other models; we observe that SceneSayerSDE is up to $\sim 70\%$ better on the R@10 metric than the best baseline variant. Additionally, the SceneSayerODE/SDE models exhibit reduced prediction bias, as evidenced by their consistently better performance across the mean recall metrics.
@@ 6.2 Results
0.3 STTran+ [7] 22.3 22.9 22.9 25.5 38.3 45.6 8.6 9.1 9.1 13.1 24.8 42.3
DSGDetr+ [9] 13.6 14.1 14.1 22.2 33.9 46.0 4.8 5.0 5.0 7.9 16.9 40.9
STTran++ [7] 22.1 22.8 22.8 28.1 39.0 45.2 9.2 9.8 9.8 17.7 30.6 42.0
DSGDetr++ [9] 18.2 18.8 18.8 27.7 39.2 47.3 8.9 9.4 9.4 15.3 26.6 44.0
@@
0.5 STTran+ [7] 24.2 24.9 24.9 27.5 41.6 49.9 9.3 9.9 9.9 13.8 26.7 43.1
DSGDetr+ [9] 15.8 16.2 16.2 24.7 38.2 52.4 5.5 5.7 5.7 8.7 18.5 41.5
STTran++ [7] 24.5 25.2 25.2 30.6 43.2 50.2 10.1 10.7 10.7 18.4 29.5 43.1
DSGDetr++ [9] 20.7 21.4 21.4 30.4 44.0 52.7 10.2 10.8 10.8 16.5 30.8 45.1
@@
0.7 STTran+ [7] 28.9 29.4 29.4 33.6 49.9 58.7 10.7 11.3 11.3 16.3 32.2 49.9
DSGDetr+ [9] 18.8 19.1 19.1 29.9 44.7 60.2 6.5 6.8 6.8 10.4 22.4 47.7
STTran++ [7] 29.1 29.7 29.7 36.8 51.6 58.7 11.5 12.1 12.1 21.2 34.6 49.0
DSGDetr++ [9] 24.6 25.2 25.2 36.7 51.8 60.6 12.0 12.6 12.6 19.7 36.4 50.6
@@
0.9 STTran+ [7] 30.9 31.3 31.3 39.9 56.6 63.8 11.5 11.9 11.9 20.3 37.7 54.3
DSGDetr+ [9] 21.3 21.6 21.6 38.4 54.9 68.7 7.5 7.7 7.7 13.7 29.0 55.4
STTran++ [7] 31.1 31.6 31.6 43.5 57.6 63.9 12.4 12.8 12.8 25.3 39.6 54.0
DSGDetr++ [9] 27.6 28.1 28.1 45.8 61.5 68.5 13.2 13.7 13.7 25.8 42.9 58.3
@@
Partially Grounded Action Genome Scenes. We present the results in Table. 2. We observed that the STTran adaptation demonstrates superior performance compared to the DSGDetr adaptation across both model variants. Notably, the SceneSayer models significantly outperform all proposed baseline models. In particular, the SceneSayerSDE model achieves up to $\sim 30\%$ improvement on the R@10 metric compared to the best-performing baseline model. Additionally, both SceneSayerODE/SDE models outperform all proposed baseline models in mean recall metrics, suggesting their proficiency in predicting relationships with a reduced bias.
@@
0.3 STTran+ [7] 30.8 32.8 32.8 30.6 47.3 62.8 7.1 7.8 7.8 9.5 19.6 45.7
DSGDetr+ [9] 27.0 28.9 28.9 30.5 45.1 62.8 6.7 7.4 7.4 9.5 19.4 45.2
STTran++ [7] 30.7 33.1 33.1 35.9 51.7 64.1 11.8 13.3 13.3 16.5 29.3 50.2
DSGDetr++ [9] 25.7 28.2 28.2 36.1 50.7 64.0 11.1 12.8 12.8 19.7 32.0 51.1
@@
0.5 STTran+ [7] 35.0 37.1 37.1 34.4 53.4 70.8 8.0 8.7 8.8 10.5 21.5 48.8
DSGDetr+ [9] 31.2 33.3 33.3 34.3 51.0 70.8 7.8 8.6 8.6 10.5 21.4 48.4
STTran++ [7] 35.6 38.1 38.1 40.3 58.4 72.2 13.4 15.2 15.2 17.8 32.5 53.7
DSGDetr++ [9] 29.3 31.9 32.0 40.3 56.9 72.0 12.2 13.8 13.9 20.6 34.3 54.0
@@
0.7 STTran+ [7] 40.0 41.8 41.8 41.0 62.5 80.4 9.1 9.8 9.8 12.6 26.3 57.5
DSGDetr+ [9] 35.5 37.3 37.3 41.0 59.8 80.7 8.9 9.6 9.6 12.6 26.2 57.4
STTran++ [7] 41.3 43.6 43.6 48.2 68.8 82.0 16.3 18.2 18.2 22.3 39.5 63.1
DSGDetr++ [9] 33.9 36.3 36.3 48.0 66.7 81.9 14.2 15.9 15.9 24.5 41.1 63.4
@@
0.9 STTran+ [7] 44.7 45.9 45.9 50.9 74.8 90.9 10.3 10.8 10.8 16.3 33.7 71.4
DSGDetr+ [9] 38.8 40.0 40.0 51.0 71.7 91.2 10.2 10.7 10.7 16.3 33.7 71.6
STTran++ [7] 46.0 47.7 47.7 60.2 81.5 92.3 19.6 21.4 21.4 29.6 49.1 76.4
DSGDetr++ [9] 38.1 39.8 39.8 58.8 78.8 92.2 16.3 17.7 17.7 30.7 50.3 77.2
@@
Grounded Action Genome Scenes. We present the results in Table. 3. We observe that in our proposed baseline models, the adaptation of STTran surpasses the adaptation of DSGDetr for both model variations. SceneSayerODE/SDE consistently outperforms all baselines across all metrics. The SceneSayerSDE model, in particular, demonstrated remarkable efficiency, marking approximately a 26% improvement in the Recall@10 (R@10) metric compared to the previously best-performing baseline. Furthermore, both variations of the SceneSayer models, ODE and SDE, demonstrated superior performance over the conventional baseline models on mean recall metrics.
@@
Qualitative Results. In Fig. 4, we provide qualitative SGA results on AG dataset under different settings. These results are the output of baseline variants and SceneSayer on 70% observation. The comparative analysis demonstrates SceneSayer’s superior performance relative to the baseline variants. Specifically, under the AGS setting, SceneSayerODE outshines all with the fewest incorrect anticipations, closely followed by SceneSayerSDE. In the PGAGS and GAGS, SceneSayerSDE surpasses the performance of all other models.
7 Conclusion

### v2

@@ 6.1 Baselines
We select two methods from VidSGG literature, STTran [cong_et_al_sttran_2021] and DSGDetr [Feng_2021], as our strong baselines [^7] Although VidSGG literature witnessed many approaches recently, most use the representation processing pipeline proposed by the two selected transformer-based VidSGG methods.for adaptations. In our adaptation of selected methods to anticipate the relationships, we retain the object representation processing unit (ORPU) and spatial context processing unit (SCPU) as proposed originally and introduce two variants of latent dynamics processing unit (LDPU) [^8] In SceneSayer, we developed a novel ORPU but adapted the SCPU from [Feng_2021] and trained all units end-to-end from scratch. We noticed a drop in performance when we used pre-trained ORPU and SCPU (kept frozen during further training) units from [Feng_2021, cong_et_al_sttran_2021] and trained the LDPU. (see Fig. 3). The two proposed variants differ in (1) the architecture of the model and (2) the loss functions employed to train them.
In terms of architectural differences, in our baseline+ variants, we employ an anticipatory transformer built using the vanilla transformer architecture [vaswani_et_al_2017] to generate future relationship representations by processing the relationship representations in the observed temporal context. Meanwhile, in our baseline++ variants, we introduce an additional temporal encoder to process the representations before we feed them through the anticipatory transformer. This component further refines the representations by enhancing the model’s understanding of spatio-temporal dynamics. In the context of loss functions, the approach followed by our baseline+ variants focuses solely on decoding the anticipated relationship representations. However, in contrast, our baseline++ variants adopt a more comprehensive strategy. These not only decode the anticipated representations but also simultaneously decode the observed representations. This dual-decoding approach allows for a more nuanced understanding of relationships.
6.2 Results
Action Genome Scenes. We present the results in Table. 1 where we observe that both STTran++ and DSGDetr++ consistently outperform their basic counterparts. Furthermore, it can be clearly noticed that the SceneSayerSDE model consistently outperforms the SceneSayerODE model, which, in turn, performs better than the baseline variants. Although SceneSayerODE/SDE models fall short on the R@50 metric in the No Constraint graph generation strategy we note that the R@K metric holds importance for lower values of K and, in R@10 and R@20 metrics, they consistently outperform baselines. We note that SceneSayerSDE is up to $\sim 70\%$ better on the R@10 metric than the best baseline variant. Additionally, the SceneSayerODE/SDE models exhibit reduced prediction bias, as evidenced by their performance across the mean recall metrics.
@@ 6.2 Results
0.3 STTran+ [cong_et_al_sttran_2021] 12.5 19.6 20.7 13.9 21.6 40.8 3.4 5.3 5.7 3.5 7.3 20.3
DSGDetr+ [Feng_2021] 12.8 19.6 20.4 14.3 21.8 41.3 3.5 5.3 5.6 3.6 7.6 21.2
STTran++ [cong_et_al_sttran_2021] 18.5 27.9 29.5 15.4 27.2 48.6 5.9 10.4 11.3 6.2 14.1 31.2
DSGDetr++ [Feng_2021] 19.5 28.3 29.4 16.8 29.0 48.9 6.0 10.3 11.0 8.4 16.7 32.3
@@
0.5 STTran+ [cong_et_al_sttran_2021] 13.1 21.1 22.2 14.9 22.6 42.9 3.6 5.8 6.2 3.7 7.6 21.4
DSGDetr+ [Feng_2021] 13.6 20.9 21.9 15.2 23.1 43.3 3.8 5.8 6.1 3.9 8.0 22.2
STTran++ [cong_et_al_sttran_2021] 19.7 30.2 31.8 16.6 29.1 51.5 6.3 11.3 12.3 6.6 14.7 33.4
DSGDetr++ [Feng_2021] 20.7 30.3 31.6 17.4 30.5 51.9 6.4 11.0 11.7 8.4 17.0 33.9
@@
0.7 STTran+ [cong_et_al_sttran_2021] 14.9 23.4 24.7 16.6 25.1 47.2 4.1 6.5 7.0 4.2 8.5 24.0
DSGDetr+ [Feng_2021] 15.5 23.4 24.3 16.8 25.3 47.4 4.3 6.5 6.9 4.3 8.8 24.7
STTran++ [cong_et_al_sttran_2021] 22.1 33.6 35.2 19.0 32.8 56.8 7.0 12.6 13.6 7.7 17.1 36.8
DSGDetr++ [Feng_2021] 22.9 33.6 34.9 19.8 34.1 56.7 7.1 12.6 13.3 9.5 19.2 37.2
@@
0.9 STTran+ [cong_et_al_sttran_2021] 15.2 24.1 25.4 17.5 26.8 49.6 4.5 7.0 7.5 4.6 9.2 24.3
DSGDetr+ [Feng_2021] 16.2 24.8 25.9 17.9 27.7 51.4 4.8 7.2 7.6 4.7 9.7 25.9
STTran++ [cong_et_al_sttran_2021] 23.6 35.5 37.4 20.2 35.0 60.2 7.4 13.4 14.6 8.9 18.4 38.8
DSGDetr++ [Feng_2021] 24.4 36.1 37.6 22.2 37.1 61.0 7.4 13.8 14.8 11.4 21.0 39.5
@@
0.3 STTran+ [cong_et_al_sttran_2021] 22.3 22.9 22.9 25.5 38.3 45.6 8.6 9.1 9.1 13.1 24.8 42.3
DSGDetr+ [Feng_2021] 13.6 14.1 14.1 22.2 33.9 46.0 4.8 5.0 5.0 7.9 16.9 40.9
STTran++ [cong_et_al_sttran_2021] 22.1 22.8 22.8 28.1 39.0 45.2 9.2 9.8 9.8 17.7 30.6 42.0
DSGDetr++ [Feng_2021] 18.2 18.8 18.8 27.7 39.2 47.3 8.9 9.4 9.4 15.3 26.6 44.0
@@
0.5 STTran+ [cong_et_al_sttran_2021] 24.2 24.9 24.9 27.5 41.6 49.9 9.3 9.9 9.9 13.8 26.7 43.1
DSGDetr+ [Feng_2021] 15.8 16.2 16.2 24.7 38.2 52.4 5.5 5.7 5.7 8.7 18.5 41.5
STTran++ [cong_et_al_sttran_2021] 24.5 25.2 25.2 30.6 43.2 50.2 10.1 10.7 10.7 18.4 29.5 43.1
DSGDetr++ [Feng_2021] 20.7 21.4 21.4 30.4 44.0 52.7 10.2 10.8 10.8 16.5 30.8 45.1
@@
0.7 STTran+ [cong_et_al_sttran_2021] 28.9 29.4 29.4 33.6 49.9 58.7 10.7 11.3 11.3 16.3 32.2 49.9
DSGDetr+ [Feng_2021] 18.8 19.1 19.1 29.9 44.7 60.2 6.5 6.8 6.8 10.4 22.4 47.7
STTran++ [cong_et_al_sttran_2021] 29.1 29.7 29.7 36.8 51.6 58.7 11.5 12.1 12.1 21.2 34.6 49.0
DSGDetr++ [Feng_2021] 24.6 25.2 25.2 36.7 51.8 60.6 12.0 12.6 12.6 19.7 36.4 50.6
@@
0.9 STTran+ [cong_et_al_sttran_2021] 30.9 31.3 31.3 39.9 56.6 63.8 11.5 11.9 11.9 20.3 37.7 54.3
DSGDetr+ [Feng_2021] 21.3 21.6 21.6 38.4 54.9 68.7 7.5 7.7 7.7 13.7 29.0 55.4
STTran++ [cong_et_al_sttran_2021] 31.1 31.6 31.6 43.5 57.6 63.9 12.4 12.8 12.8 25.3 39.6 54.0
DSGDetr++ [Feng_2021] 27.6 28.1 28.1 45.8 61.5 68.5 13.2 13.7 13.7 25.8 42.9 58.3
@@
Partially Grounded Action Genome Scenes. We present the results in Table. 2. We observed that the STTran adaptation demonstrates superior performance compared to the DSGDetr adaptation across both model variants. Notably, the SceneSayer models significantly outperform all proposed baseline models. In particular, the SceneSayerSDE model achieves up to $\sim 30\%$ improvement on the R@10 metric compared to the best-performing baseline model and both SceneSayerODE/SDE models outperform all proposed baseline models in mean recall metrics.
Grounded Action Genome Scenes. We present the results in Table. 3 and we observe that in our proposed baseline models, the adaptations of STTran perform better than the adaptations of DSGDetr and SceneSayer variants outperformed all baselines.
@@
0.3 STTran+ [cong_et_al_sttran_2021] 30.8 32.8 32.8 30.6 47.3 62.8 7.1 7.8 7.8 9.5 19.6 45.7
DSGDetr+ [Feng_2021] 27.0 28.9 28.9 30.5 45.1 62.8 6.7 7.4 7.4 9.5 19.4 45.2
STTran++ [cong_et_al_sttran_2021] 30.7 33.1 33.1 35.9 51.7 64.1 11.8 13.3 13.3 16.5 29.3 50.2
DSGDetr++ [Feng_2021] 25.7 28.2 28.2 36.1 50.7 64.0 11.1 12.8 12.8 19.7 32.0 51.1
@@
0.5 STTran+ [cong_et_al_sttran_2021] 35.0 37.1 37.1 34.4 53.4 70.8 8.0 8.7 8.8 10.5 21.5 48.8
DSGDetr+ [Feng_2021] 31.2 33.3 33.3 34.3 51.0 70.8 7.8 8.6 8.6 10.5 21.4 48.4
STTran++ [cong_et_al_sttran_2021] 35.6 38.1 38.1 40.3 58.4 72.2 13.4 15.2 15.2 17.8 32.5 53.7
DSGDetr++ [Feng_2021] 29.3 31.9 32.0 40.3 56.9 72.0 12.2 13.8 13.9 20.6 34.3 54.0
@@
0.7 STTran+ [cong_et_al_sttran_2021] 40.0 41.8 41.8 41.0 62.5 80.4 9.1 9.8 9.8 12.6 26.3 57.5
DSGDetr+ [Feng_2021] 35.5 37.3 37.3 41.0 59.8 80.7 8.9 9.6 9.6 12.6 26.2 57.4
STTran++ [cong_et_al_sttran_2021] 41.3 43.6 43.6 48.2 68.8 82.0 16.3 18.2 18.2 22.3 39.5 63.1
DSGDetr++ [Feng_2021] 33.9 36.3 36.3 48.0 66.7 81.9 14.2 15.9 15.9 24.5 41.1 63.4
@@
0.9 STTran+ [cong_et_al_sttran_2021] 44.7 45.9 45.9 50.9 74.8 90.9 10.3 10.8 10.8 16.3 33.7 71.4
DSGDetr+ [Feng_2021] 38.8 40.0 40.0 51.0 71.7 91.2 10.2 10.7 10.7 16.3 33.7 71.6
STTran++ [cong_et_al_sttran_2021] 46.0 47.7 47.7 60.2 81.5 92.3 19.6 21.4 21.4 29.6 49.1 76.4
DSGDetr++ [Feng_2021] 38.1 39.8 39.8 58.8 78.8 92.2 16.3 17.7 17.7 30.7 50.3 77.2
@@
Specifically, the proposed SceneSayerSDE model demonstrated remarkable efficiency, marking approximately a 26% improvement in the R@10 metric compared to the best-performing baseline. Furthermore, both SceneSayerSDE/ODE variants demonstrated superior performance consistently over the baseline variants on the mean recall metrics.
Qualitative Results. In Fig. 7, we provide qualitative SGA results on AG dataset under different settings. These results are the output of baseline variants and SceneSayer on 70% observation. The comparative analysis demonstrates SceneSayer’s superior performance. Specifically, under the AGS setting, SceneSayerODE outshines all with the fewest incorrect anticipations, closely followed by SceneSayerSDE. In the PGAGS and GAGS, SceneSayerSDE surpasses the performance of all other models.
@@
7 Discussion, Limitations & Future Work